# Schema-Driven Information Extraction from Heterogeneous Tables

### v3

@@ 4.6 Knowledge Distillation
Considering the strong performance of API-based models on Schema-Driven IE, we now show that it is possible to use knowledge distillation (Le et al., 2022; Kang et al., 2023) to build a cost-efficient compact model, using ML tables as a demonstration. Specifically, this process first generates synthetic data by performing inference on unlabeled tables using code-davinci-002, followed by fine-tuning a smaller model (e.g., 7B parameters) using the synthetic data. We compile a collection of 979 arXiv ML papers, submitted between 2008 and 2019, yielding 3,434 tables (containing a total of 100K cells). In Table 3, we can see that LLaMA-7B and Alpaca-7B demonstrate similar performance as seen in the fine-tuning results (Table 2). While fine-tuning LLaMA with LoRA (Hu et al., 2022) presents noticeable computational efficiency, full-parameter fine-tuning of T5-11B matches the teacher model’s performance. [^12] As we observe that the T5-11B student model slightly outperforms the teacher model, we conduct a statistical significance test following Berg-Kirkpatrick et al. (2012). With 1000 bootstrap samples, the p-value is 42.3%, suggesting that the performance gap between the T5-11B student model and the teacher model is not statistically significant.
@@
4.7 Positive Externalities of InstrucTE
To further validate InstrucTE’s practicality, we integrate it with multi-modal models, like GPT4-V, for extracting data from table images. In an initial study with ML tables, it yields a Table-F1 of 70.2, approaching the 74.2 Table-F1 achieved with the original text inputs. Additionally, we explore InstrucTE’s application to the task of Leaderboard Extraction, where it shows competitive performance against leading supervised systems. Due to space constraints, details on these explorations are provided in Appendix F.
@@ 5 Related Work
Recently there have been many research efforts involving tables, particularly, table-to-text generation (Parikh et al., 2020; Wang et al., 2022; Hu et al., 2023). For example, ToTTo Parikh et al. (2020) introduced the task of open-domain table-to-text generation. Another burgeoning area is question answering and fact verification on tables (Jauhar et al., 2016; Zhong et al., 2017; Yu et al., 2018; Schlichtkrull et al., 2021). Various approaches have emerged, such as semantic parsing for compositional question answering (Pasupat and Liang, 2015), and symbolic reasoning for fact verification (Chen et al., 2020). In contrast, our work transforms tables into structured JSON records, where a data schema is the only supervision provided.
@@
The rise of pre-trained language models Devlin et al. (2019), has stimulated interest in pre-training on semi-structured data. TaPas Herzig et al. (2020) and TaBERT Yin et al. (2020) pre-train on linearized tables with a specialized cell index embedding. TABBIE Iida et al. (2021) employs dual transformers for separate row and column encoding. HTLM Aghajanyan et al. (2022) uses an HTML-specialized pre-training objective, facilitating a novel structured prompting scheme. Similar to our work, TabLLM (Hegselmann et al., 2023) uses general-purpose LLMs to process linearized tables, but we focus on schema-driven IE rather than table classification or question answering.
@@
Information extraction from semi-structured data has gained increasing interest (Carlson and Schafer, 2008; Dong et al., 2020; Gupta et al., 2022; Lou et al., 2023). Innovations such as OpenCeres (Lockard et al., 2019) and ZeroShotCeres (Lockard et al., 2020) highlight open-domain extraction from web data, while AxCell (Kardas et al., 2020) and TDMS-IE (Hou et al., 2019) focus on leaderboard extraction from ML tables. DisCoMat (Gupta et al., 2022) showcases material composition extraction from scientific tables. Despite these advancements, most existing methods require supervised datasets or specialized models for task-specific fine-tuning. Our approach stands out by using LLMs to accurately extract data across various formats and domains without relying on labels or custom extraction pipelines.
@@ 6 Conclusion
This paper explores the capabilities of LLMs for extracting structured data from heterogeneous tables. We introduce a new task, Schema-Driven Information Extraction, which converts tables into structured records guided by a human-authored data schema. To facilitate this task, we present a benchmark, comprised of tables from four diverse domains, and evaluate various LLMs through our proposed method InstrucTE. The experiments reveal that while API-based models excel across domains and formats, open-source models display significant potential in specific areas. Moreover, we conduct detailed ablation studies and analyses to investigate the factors contributing to model success, and validate the feasibility of building more compact models through distillation to reduce dependency on APIs. These contributions lay a solid foundation for the continued advancement of efficient open-source table data extraction.
@@
While InstrucTE showcases strong performance as an instruction-based prompting approach, it encounters specific challenges. Firstly, its effectiveness is highly dependent on how prompts are phrased. Despite offering guidelines for crafting prompts in Appendix A, e.g., emphasizing the use of clear attribute names and the selection of simpler attributes, a systematic method for developing robust extraction schemas for new domains remains elusive, often relying on iterative experimentation. Additionally, the model’s varying performance across different domains and formats is difficult to interpret, possibly due to biases in the pretraining corpus, a factor we cannot fully analyze due to the opaque nature of the pre-training process. InstrucTE also faces difficulties with dataset-specific nuances, as it operates on general task descriptions without detailed examples, making it challenging to navigate boundary cases effectively.
@@ Acknowledgments
We would like to thank Azure’s Accelerate Foundation Models Research Program and OpenAI’s Researcher Access Program for graciously providing access to API-based models, such as GPT-4. This research is supported in part by the NSF (IIS-2052498), ODNI and IARPA via the HIATUS program (2022-22072200004), and the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001119C0108. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. This work is approved for Public Release, Distribution Unlimited.
@@ Appendix A InstrucTE
ML Tables
[Retrieve paragraphs]
[Input table]
Here are JSON templates for four types of numeric cells: “Other”, “Result”, “Data Stat.”, and “Hyper-parameter/Architecture”:
{“value”: “xx”, “type”: “Other”}
{“value”: “xx”, “type”: “Result”, “task”: “xx”, “metric”: “xx”, “training data/set”: “xx”, “test data/set”: “xx”, “model/method”: “xx”, “model/method settings”: {“xx”: “yy”}, “experimental settings”: {“xx”: “yy”}}
{“value”: “xx”, “type”: “Data Stat.”, “dataset”: “xx”, “attribute name”: “xx”, “sub-set/group name”: “xx”, “dataset features”: {“xx”: “yy”}}
{“value”: “xx”, “type”: “Hyper-parameter/Architecture”, “model”: “xx”, “parameter/architecture name”: “xx”, “dataset”: “xx”}
Please describe all numeric cells in the above latex table following the JSON templates (proceeding by row in a left-right, top-down direction). For each cell, output one JSON description per line. For any unanswerable attributes in the templates, set their value to the placeholder “xx” if it is of string type and {“xx”: “yy”} if it is of dictionary type.
Cell Description:
{“value”: “[Query cell]”, “type”:
Chem. Tables
[Input table]
Here are JSON templates for six types of numeric cells: “Other”, “IC50”, “EC50”, “CC50”, “MIC”, and “GI50”:
{“value”: “xx”, “type”: “Other”}
{“value”: “xx”, “type”: “IC50”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
{“value”: “xx”, “type”: “EC50”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
{“value”: “xx”, “type”: “CC50”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
{“value”: “xx”, “type”: “MIC”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
{“value”: “xx”, “type”: “GI50”, “unit”: “xx”, “treatment compound”: “xx”, “target compound”: “xx”}
Please describe all numeric cells in the above XML table following the JSON templates (proceeding by row in a left-right, top-down direction). For each cell, output one JSON description per line. For any unanswerable attributes in the templates, set their value to the placeholder “xx”.
Cell Description:
{“value”: “[Query cell]”, “type”:
@@
DisCoMat
[Input table]
Here are JSON templates for two types of numeric cells: “Other” and “Glass_Compound_Amount”:
{“value”: “xx”, “type”: “Other”}
{“value”: “xx”, “type”: “Glass_Compound_Amount”, “constituent compound name”: “xx”, “unit”: “xx”, “glass material/sample name/id/code”: “xx”}
Please describe all numeric cells in the above table following the JSON templates (proceeding by row in a left-right, top-down direction). For each cell, output one JSON description per line. For any unanswerable attributes in the templates, set their value to the placeholder “xx”.
Cell Description:
{“value”: “[Query cell]”, “type”:
SWDE-auto
[Input webpage]
Here is the JSON template for automobile attribute extraction:
{“webpage title”: “xx”, “automobile model (year)”: “xx”, “price”: “xx”, “engine type”: “xx”, “fuel economy”: “xx”}
Please extract the automobile’s attributes from the HTML code above following the JSON template. For any unanswerable attributes in the template, set their value to the placeholder “<NULL>”.
{“webpage title”: “[webpage title]”, “automobile model (year)”:
@@ Appendix E Error Analysis of Caption
Appendix F Positive Externalities of InstrucTE
F.1 Extraction from Table Images
@@
F.2 Leaderboard Extraction from ML Papers

### v4

@@ 4.6 Knowledge Distillation
Considering the strong performance of API-based models on Schema-Driven IE, we now show that it is possible to use knowledge distillation (Le et al., 2022; Kang et al., 2023) to build a cost-efficient compact model, using ML tables as a demonstration. Specifically, this process first generates synthetic data by performing inference on unlabeled tables using code-davinci-002, followed by fine-tuning a smaller model (e.g., 7B parameters) using the synthetic data. We compile a collection of 979 arXiv ML papers, submitted between 2008 and 2019, yielding 3,434 tables (containing a total of 100K cells). In Table 4, we can see that LLaMA-7B and Alpaca-7B demonstrate similar performance as seen in the fine-tuning results (Table 2). While fine-tuning LLaMA with LoRA (Hu et al., 2022) presents noticeable computational efficiency, full-parameter fine-tuning of T5-11B matches the teacher model’s performance.131313 The improvement over the teacher model is not significant (p-value is 42.3%, Berg-Kirkpatrick et al., 2012).
@@
4.7 Leaderboards and Image Extraction
To further validate InstrucTE’s practicality, we integrate it with multi-modal models, like GPT4-V, for extracting data from table images. In an initial study with ML tables, it yields a Table-F1 of 70.2, approaching the 74.2 Table-F1 achieved with the original text inputs. Additionally, we explore InstrucTE’s application to the task of Leaderboard Extraction, where it shows competitive performance against leading supervised systems. Due to space constraints, details on these explorations are provided in Appendix G.
@@ 5 Related Work
Recently there have been many research efforts involving tables, particularly, table-to-text generation (Parikh et al., 2020; Wang et al., 2022; Hu et al., 2023). For example, ToTTo Parikh et al. (2020) introduced the task of open-domain table-to-text generation. In contrast, our work transforms tables into structured JSON records, where a data schema is the only supervision provided.
@@
TaPas Herzig et al. (2020) and TaBERT Yin et al. (2020) pre-train on linearized tables with a specialized cell index embedding. TABBIE Iida et al. (2021) employs dual transformers for separate row and column encoding. Similarly, TabLLM (Hegselmann et al., 2023) uses general-purpose LLMs to process tables, but we focus on schema-driven IE rather than table classification or question answering.
@@
Information extraction from semi-structured data has gained increasing interest (Carlson and Schafer, 2008; Dong et al., 2020; Gupta et al., 2022; Lou et al., 2023). OpenCeres (Lockard et al., 2019) and ZeroShotCeres (Lockard et al., 2020) highlight open-domain extraction from web data, while AxCell (Kardas et al., 2020) and TDMS-IE (Hou et al., 2019) focus on leaderboard extraction from ML tables. DisCoMat (Gupta et al., 2022) showcases material composition extraction from scientific tables. Unlike most existing methods requiring supervised datasets for fine-tuning, our approach stands out by using LLMs to accurately extract data across various domains using an extraction schema.
@@ 6 Conclusion
This paper explores the capabilities of LLMs for extracting structured data from heterogeneous tables. We introduce a new task, Schema-Driven Information Extraction, which converts tables into structured records guided by a human-authored data schema. To facilitate this task, we present a benchmark, comprised of tables from four diverse domains, and evaluate various LLMs through our proposed method InstrucTE. The experiments reveal that while API-based models excel across domains and formats, open-source models display significant potential in specific areas. Moreover, we conduct detailed ablation studies and analyses to investigate the factors for model success, and validate the feasibility of building compact models through distillation to reduce dependency on APIs.
@@
While InstrucTE showcases strong performance as an instruction-based prompting approach, it encounters specific challenges. Firstly, similar to other prompting methods, its performance could be sensitive to the phrasing of the prompt. Despite offering guidelines for crafting prompts in Appendix A, such as emphasizing clear attribute names, developing robust extraction schemas for new domains often relies on iterative experimentation. Future work could explore automatic prompt optimization (Zhou et al., 2023; Wang et al., 2023) to reduce the need for human trial-and-error. Additionally, the model’s varying performance across different domains and formats is difficult to interpret, possibly due to biases in the pretraining corpus, a factor we cannot fully analyze due to the opaque nature of the pre-training process. InstrucTE also faces difficulties with dataset-specific nuances, as it operates on general task descriptions without detailed examples, making it challenging to navigate boundary cases effectively.
@@ Acknowledgments
We would like to thank Azure’s Accelerate Foundation Models Research Program and OpenAI’s Researcher Access Program for graciously providing access to API-based models, such as GPT-4. This research is supported in part by the NSF (IIS-2052498), ODNI and IARPA via the HIATUS program (2022-22072200004), and the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001119C0108. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. This work is approved for Public Release, Distribution
@@ Appendix A InstrucTE
ML Tables [Uncaptioned image]
Chem. Tables [Uncaptioned image]
@@
DisCoMat [Uncaptioned image]
SWDE-auto [Uncaptioned image]
@@
Dataset # Records # Attr. Precision IAA
ML Tables 100 591 73.9 95.7
Chem. Tables 100 380 95.3 100
DisCoMat 100 201 92.5 99.4
SWDE 160 640 96.4 98.2
Table 9: Statistics and results of attribute-level human evaluation on four datasets. The inter-annotator agreement score (IAA) is calculated as the F1 score between the two annotations
@@ Appendix E Error Analysis of Caption
Appendix F InstrucTE Errors
This section presents examples of InstrucTE’s errors (w/ GPT-4) on ML tables, illustrating each of the eight fine-grained error types.
F.1 False Negative: gold answer in table, not predicted
Input (Table 5 of arXiv paper 2210.00044v1)
Predicted Records
1{"value": "83.48", ..., "model/method settings": {"xx": "yy"}, ...}
2{"value": "62.81", ..., "model/method settings": {"xx": "yy"}, ...}
3 ...
4{"value": "94.32", ..., "model/method settings": {"xx": "yy"}, ...}
5 ...
Gold Records
1{"value": "83.48", ..., "model/method settings": {"w/o Pretraining": "true"}, ...}
2{"value": "62.81", ..., "model/method settings":{"w/o Pretraining": "true"}, ...}
3 ...
4{"value": "94.32", ..., "model/method settings": {"w Pretraining": "true"}, ...}
5 ...
F.2 False Negative: gold answer in main text, not predicted
Input (Table 4 of arXiv paper 2210.00193v1)
[Uncaptioned image][Uncaptioned image]
Predicted Records
1{"value": "51.2", ..., "training data/set": "xx", ...}
2{"value": "50.9", ..., "training data/set": "xx", ...}
3 ...
4{"value": "66.7", ..., "training data/set": "xx", ...}
5 ...
Gold Records
1{"value": "51.2", ..., "training data/set": "a mixture of monolingual and parallel data from 112 languages mined from the web", ...}
2{"value": "50.9", ..., "training data/set": "a mixture of monolingual and parallel data from 112 languages mined from the web", ...}
3 ...
4{"value": "66.7", ..., "training data/set": "a mixture of monolingual and parallel data from 112 languages mined from the web", ...}
5 ...
F.3 False Negative: gold answer predicted, wrong attribute
Input (Table 5 of arXiv paper 2210.00044v1)
Predicted Records
1{"value": "83.48", ..., "model": "EWC", "experimental settings": {"xx": "yy"}, ...}
2{"value": "62.81", ..., "model": "EWC", "experimental settings": {"xx": "yy"}, ...}
3 ...
4{"value": "94.32", ..., "model": "EWC", "experimental settings": {"xx": "yy"}, ...}
5 ...
Gold Records
1{"value": "83.48", ..., "experimental settings": {"Method": "EWC"}, ...}
2{"value": "62.81", ..., "experimental settings": {"Method": "EWC"}, ...}
3 ...
4{"value": "94.32", ..., "experimental settings": {"Method": "EWC"}, ...}
5 ...
F.4 False Positive: gold answer in main text, table header predicted
Input (Table 5 of arXiv paper 2210.00044v1)
[Uncaptioned image][Uncaptioned image]
Predicted Records
1{"value": "83.48", ..., "training data/set": "What room", ...}
2{"value": "62.81", ..., "training data/set": "What sport", ...}
3 ...
4{"value": "94.32", ..., "training data/set": "What room", ...}
5 ...
Gold Records
1{"value": "83.48", ..., "training data/set": "VQA-v2", ...}
2{"value": "62.81", ..., "training data/set": "VQA-v2", ...}
3 ...
4{"value": "94.32", ..., "training data/set": "VQA-v2", ...}
5 ...
F.5 False Positive: partial match, but misses important details
Input (Table 4 of arXiv paper 2210.00193v1)
Predicted Records
1{"value": "51.2", ..., "test data/set": "FRMT test", ...}
2{"value": "50.9", ..., "test data/set": "FRMT test", ...}
3 ...
4{"value": "66.7", ..., "test data/set": "FRMT test", ...}
5 ...
Gold Records
1{"value": "51.2", ..., "test data/set": "FRMT pt", ...}
2{"value": "50.9", ..., "test data/set": "FRMT zh", ...}
3 ...
4{"value": "66.7", ..., "test data/set": "FRMT pt", ...}
5 ...
F.6 False Positive: gold answer in table caption, table header predicted
Input (Table 3 of arXiv paper 2210.00740v1)
Predicted Records
1{"value": "97.2", ..., "metric": "Hea", ...}
2{"value": "86.7", ..., "metric": "Wri", ...}
3 ...
4{"value": "82.3", ..., "metric: "Ank", ...}
5 ...
Gold Records
1{"value": "97.2", ..., "metric": "AP", ...}
2{"value": "86.7", ..., "metric": "AP", ...}
3 ...
4{"value": "82.3", ..., "metric": "AP", ...}
5 ...
F.7 False Positive: complete mismatch
Input (Table 3 of arXiv paper 2210.00740v1)
Predicted Records
1{"value": "97.2", ..., "experimental settings": {"Venue": "CVPR 2020"}, ...}
2{"value": "86.7", ..., "experimental settings": {"Venue": "CVPR 2020"}, ...}
3 ...
4{"value": "82.3", ..., "experimental settings": {"Venue": "ICCV 2021"}, ...}
5 ...
Gold Records
1{"value": "97.2", ..., "experimental settings": {"xx": "yy"}, ...}
2{"value": "86.7", ..., "experimental settings": {"xx": "yy"}, ...}
3 ...
4{"value": "82.3", ..., "experimental settings": {"xx": "yy"}, ...}
5 ...
F.8 Propagated Errors: select a wrong record template
Input (Table 7 of arXiv paper 2210.00627v1)
Predicted Records
1{"value": "3", "type": "Hyper-parameter", ...}
2{"value": "6", "type": "Hyper-parameter", ...}
3 ...
4{"value": "14", "type": "Hyper-parameter", ...}
Gold Records
1{"value": "3", "type": "Other"}
2{"value": "6", "type": "Other"}
3 ...
4{"value": "14", "type": "Other"}
Appendix G Extracting Leaderboards from Table Images
G.1 Extraction from Table Images
@@
G.2 Leaderboard Extraction from ML Papers